OpenStudy (anonymous):

find the temperature at depth 2 kilometers T(d)=35d+20 where d is depth (km)

OpenStudy (kamille):

hey,is it function? since tempreture depends how deep it is, you need to find T(2) T(2)=35*2+20

OpenStudy (anonymous):

T(d)=35d+20 when d = 2 km , then T(d)=35d+20 T(d)=35(2) + 20 T(d)= 70 + 20 T(d)= 90